List of Slides

1

– General Linux 1 – Search text files using regular expressions []

(Linux Professional Institute Certification)

a

   .~.
   /V\      by: geoffrey robertson
  // \\         geoffrey@zip.com.au
  @._.@
$Id: gl1.103.2.slides.tex,v 1.3 2003/05/29 14:10:18 geoffr Exp $

aCopyright \begin{matrix}c \\ ◯ \end{matrix} 2002 Geoffrey Robertson. Permission is granted to make and distribute verbatim copies or modified versions of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.

1.103.1 1.103.2 1.103.3 1.103.4 1.103.5 1.103.6 1.103.7

Objective

4

Key files, terms, and utilities

5

Resources of interest

6

The End

7